OpenStudy (ttp):

Every day, Marco runs for 40 minutes and covers 5 miles. How long would it take him to run 2 miles? 8 min 16 min 37 min 100 min

OpenStudy (the_fizicx99):

40/5 = ?

OpenStudy (anonymous):

(40*2 = 80. 80/5 = 16)

OpenStudy (the_fizicx99):

@Light&Happiness refrain from giving out answers. Anyway, 40/5 = 8 This means 1 mile every 8 minutes. So when he runs 2 miles he'll have run 16 minutes.

OpenStudy (nikato):

40 x min ---- = --------- 5 mi 2 mi and you can start by cross muliptli\ying and you end up with something like those^

OpenStudy (31356):

Yes! 16 is the answer. Set up a proportion! Minutes over miles or miles over minutes then multiply then divide .

OpenStudy (31356):

Medal? If helped.......

OpenStudy (the_fizicx99):

Why are you always asking for medals smh